# Great Models Think Alike and this Undermines AI Oversight

### v1

@@ Abstract
Shashwat Goel1,2∘  Joschka Strüber3,4∘  Ilze Amanda Auzina3,4∘  Karuna K Chandra5∘
Ponnurangam Kumaraguru5  Douwe Kiela6,7  Ameya Prabhu3,4∘  Matthias Bethge3,4  Jonas Geiping1,2,3
1ELLIS Institute Tübingen  2Max Planck Institute for Intelligent Systems  3Tübingen AI Center
4University of Tübingen  5IIIT Hyderabad  6Contextual AI  7Stanford University  core contributors
[…]
@@ 1 Introduction
Machine Learning model capabilities have improved immensely over the last few years. Scaling up the amount of data used for training has played a crucial role in these improvements (Kaplan et al., 2020). Initially, most of the gains in Language Model (LM) capabilities came from scaling pretraining data (Llama Team, 2024a). Recently, there is increasing interest in post-training, either with human preferences (Ouyang et al., 2022), or task-specific expert annotations (Lightman et al., 2023). Collecting human preferences or annotations is slow and expensive. Therefore, with increasing model capabilities, an attractive alternative is to use LMs to annotate training data  (Gilardi et al., 2023) and score model outputs (Zheng et al., 2023), to boost both training (Stiennon et al., 2020) and evaluation (Li et al., 2024b). In this paper, we refer to both these techniques together as AI oversight [^1] The term is inspired by “scalable oversight” (Bowman et al., 2022), which studies human-in-the-loop mechanisms for AI Safety..
Can we rely on AI oversight going forward? This remains a topic of much debate. In this work, we study oversight from the perspective of model similarity. When assessing or teaching humans, it is well recognized that individuals have different strengths and weaknesses. Similarly, two models with 50% accuracy may misclassify completely different samples and thus be highly dissimilar (having different ‘strengths’). To measure model similarity, we build on error consistency (Geirhos et al., 2020), which measures overlap in the samples where two models err beyond what is expected by chance due to the two models’ accuracies. In Section 2, we extend the error consistency metric in two crucial ways – 1) by counting differences in predictions rather than correctness for each sample, and 2) incorporating output probabilities. In this way, our novel similarity metric, Chance Adjusted Probabilistic Alignment (CAPA), provides a novel way to quantify functional similarity between models. We use this to analyze both evaluation and training using AI oversight as depicted in Figure 1:
@@ 2.2 Our Contribution
Alternatives and Justification. We summarize comparisons to existing functional similarity measures based on key desiderata ( 1- 3) in Table 1. In Appendix A.4 we justify design choices for CAPA, comparing it with various alternatives like using Jensen Shannon Distance (JSD), or defining $c_{\text{exp}}^{p}$ using assumptions similar to Scott’s $\pi$ instead of Cohen’s $\kappa$. We provide plots with alternative similarity metrics for our main empirical takeaways throughout the Appendix, and find consistent trends. In each case, $\kappa_{p}$ shows the trend most clearly, with the least noise. CAPA can also be used when probabilities are unavailable by assigning probability $1$ to the predicted option and 00 to the others. We use this to prove $\kappa_{p}$ is a strict generalization of error consistency, and reduces to it for binary classification (Appendix A.1). Furthermore, $\kappa_{p}$ can be extended beyond pairwise comparisons to multiple models, (Appendix A.2). For completeness, we present probabilistic extensions for Cohen’s $\kappa$, Scott’s $\pi$, Fleiss’ $\kappa_{F}$ in Appendix A.5 and show comparisons to CAPA on illustrative examples and synthetic data in Appendix A.7.
@@ 3 Affinity Bias in LLM-as-a-Judge
Evaluating free-response model generations automatically at scale is tricky (Biderman et al., 2024). This has led to popular leaderboards like Arena-hard-auto (Li et al., 2024b), AlpacaEval 2.0 (Dubois et al., 2024), and AidanBench (McLaughlin et al., 2024) adopting LLM-as-a-Judge for scoring. Recent work cautions that language models are biased towards their own outputs (Liu et al., 2024; Panickssery et al., 2024), and these leaderboards assume that excluding the judge model from the rankings circumvents this problem. However, it has been shown that human interviewers are biased towards candidates with similar knowledge and traits, a phenomenon called affinity bias (Bagues & Perez-Villadoniga, 2012). We study whether LMs also exhibit a bias toward similar models. This would indicate that it is not sufficient to just use a held-out LM as the judge; one must account for the confounder of similarity.
@@ 3.1 Experimental Setup
To study whether LM judges prefer more similar models, we evaluate a large set of judges and models on MMLU-Pro (Wang et al., 2024), a benchmark for hard problem solving questions across 14 domains. We filter 8,707 questions that can also be answered in a free-text response style, without options, following Myrzakhan et al. (2024). Each question is posed to every evaluated model as an MCQ and as an open-style question. The per-sample results of the former are used to compute the similarities of judge-model pairs, whereas responses to the latter are given to an LLM-as-a-judge. The judge has to make a binary decision on whether a free-text response is correct or wrong. This is done based on its own internal knowledge without access to a ground-truth solution. We call the average of binary judgments across the benchmark the model’s Judgment Score for a given judge. Using a parallel MCQ evaluation with ground-truth answers allows us to compare the judgment scores with verifiable accuracy measurements (details and comparisons to alternatives are in Appendix B.2), consistent with prior scalable oversight research (Bowman et al., 2022). We compute pairwise similarity with CAPA, $\kappa_{p}$, across 9 judge and 39 model pairs. For a complete overview of models investigated, the question filtering process, the inference setup and the prompts used for judges see Appendix B.
@@ 3.2 Results & Discussion
Figure 2: Judgment Score relation with Model Similarity. Each line is a regression model fit between judgment and similarity scores as computed between model and judge pairs. Each point represents a single pair, and $\diamond$ indicates that both, the model and the judge, come from the same model family. We report for each fit the corresponding Pearson correlation values, $r$. We found significant positive correlation between judgment scores and similarity across all judges, **∗ ∗ indicates $p<0.01$.
@@
Q1: Do LM Judges favor more similar models? As a motivating example, Qwen2.5-72B-Instruct as a judge scores Qwen2.5-7B-Instruct as being 71% correct, while the more capable (41% vs 51% MCQ accuracy) Llama-3.1-70B-Instruct is deemed less accurate at 67%. In Figure 2 we show that model favoritism extends beyond self- or family- preference to all models that are functionally similar. We find a significant ($p<0.01$) positive correlation (average Pearson r=$0.84$) between LLM-as-a-judge scores and model similarity ($\kappa_{p}$) for all judges.
Q2: Is this merely due to better accuracy? Note that while $\kappa_{p}$ adjusts for inflation in agreement of highly accurate models, we do expect models with lower accuracy to be less similar with highly capable judge models, and thus have lower $\kappa_{p}$. To control for the accuracy of the evaluated model we perform multiple regression and partial correlation analysis (see Table 2). The multiple regression analysis shows that both, accuracy and similarity, have a significant positive effect on the judge score. The coefficient of accuracy increases for more capable judge models, consistent with prior work showing improved alignment with human judges (Thakur et al., 2024). We find that especially for small models ($<$32B) the effect of similarity is greater. The partial correlation results control for accuracy and confirm that there is still a significant effect of similarity on judgment scores even for the best judge models. Altogether, the statistical analysis confirms that judgment scores are confounded by affinity bias.
@@ 4.1 Experimental Setup
We measure similarity between the weak supervisor and base student model on the validation set. We then perform weak-to-strong training on the student model, using the confidence-weighted finetuning objective proposed in Burns et al. (2024). We investigate if similarity is an apriori predictor of performance gained on the test set. For our experiments, we study 4 weak models in the $1-3$B parameter range, and 3 strong models in the $7-9$B parameter range, for a total of 12 model pairs, and $15$ of the NLP tasks studied in Burns et al. (2024), specified in Table 17. The full setup is consistent with the open-weight model reproduction by (Scherlis et al., 2024), and is described in Appendix C.1.
@@ 4.2 Results & Discussion
Q1: Does Complementary Knowledge Influence Performance Gain? Figure 3 shows that for all model combinations, similarity between the weak supervisor and initial strong student inversely correlates with the improvement obtained from weak-to-strong training ($r=-0.85)$. Even after using partial correlation analysis to control for the accuracy gap between the weak supervisor and strong student, similarity is inversely correlated with weak-to-strong gain ($r=-0.35,p<0.01$). Thus, tasks where the supervisor and student make less correlated errors tend to yield greater improvements. This contributes towards understanding why gains from weak to strong training vary across tasks, an open question posed by Burns et al. (2024).
Q2. Does Complementary Knowledge Add Beyond Elicitation? The original explanation for performance gains from weak-to-strong generalization is that the weak supervisor “elicits” the latent knowledge in the superior representations of the stronger student (Burns et al., 2024). To investigate whether complementary knowledge adds to this explanation or is subsumed within it, we first obtain the strong model with “upper-bound” elicitation by finetuning it on ground-truth annotations. We refer to this as the strong elicited model. We can then separate the test data into four parts based on whether the strong elicited and weak supervisor model were correct or wrong, measuring average accuracy of the weak-to-strong model on each part to disentangle gains from different factors. The experiment setup is discussed further in Appendix C.2.
@@
Q3. Where can weak-to-strong training improve? The strong elicited model is considered to represent upper-bound performance, but as shown in Table 3, the actual ceiling is significantly higher if complementary knowledge of the weak supervisor is fully leveraged. Interestingly, on the training set, the weak-to-strong trained model shows similar accuracy on the top-left and bottom-right quadrants as shown in Figure 12. Yet, when generalizing to unseen samples, it falls back more often to its initial priors. We hope this analysis guides future work on improving weak-to-strong training methodology, by highlighting leveraging complementary knowledge as a concrete avenue for improvement.
Figure 5: Average Similarity ($\kappa_{p}$) vs Model Capability. We split 130 LMs into 5 buckets based on their accuracy percentile. For each LM we compute its mean similarity within the bucket (across models from different developers), and plot it against model accuracy. The size of the scatter points is proportional to model size. As $\kappa_{p}$ measures overlap in mistakes, the positive correlation indicates LM mistakes are getting more correlated with increasing capabilities.
@@ 5.2 Results & Discussion
Q2. Why are model errors becoming more correlated? This is an interesting research direction in itself. We perform a preliminary analysis in Appendix D.2, summarizing key conclusions here. First, we observe only a slight increase in similarity for harder questions in our datasets, indicating difficulty is not a significant confounder for this trend. We find this trend is stronger in instruction-tuned models, and using alternative architectures like Mamba (Gu & Dao, 2023) may not be enough to increase diversity.
@@ 6 Related Work
There is increasing interest in finding differences between models for applications like visual tools for comparative analytics (Strobelt et al., 2021; Kahng et al., 2024), efficient human evaluation (Boubdir et al., 2023), comparing learning algorithms (Shah et al., 2023), identifying side-effects of API updates (Eyuboglu et al., 2024) or quantization (Dutta et al., 2024). Prior work has also looked at qualitatively describing differences between data distributions (Zhong et al., 2022, 2023; Dunlap et al., 2024b, a). Our work proposes metrics to quantify LM differences (or similarity). Huh et al. (2024) used representation similarity metrics (Kornblith et al., 2019; Bansal et al., 2021) to show convergence in visual representations and their alignment with language representations. In contrast, we show model mistakes are becoming more correlated as capabilities improve. We measure differences in input-output behaviour, which leverages sample level evaluations (Burnell et al., 2023) such as those available on OpenLLMLeaderboard (Myrzakhan et al., 2024) and HELM (Bommasani et al., 2023)Geirhos et al. (2020) proposed measuring “error consistency” between image classifiers and humans, with Geirhos et al. (2021) showing an early trend of data-rich models making more similar mistakes to humans. We enrich this metric, distinguishing between different mistakes and incorporating probabilistic information.
@@
Our results on AI judges fall in a broader line highlighting their pitfalls (Zheng et al., 2024). These include biases such as favoring verbose texts or options at certain positions (Koo et al., 2024; Ye et al., 2024). Interestingly, these biases are also sometimes found in human annotators (Chen et al., 2024). In fact, there is rich literature documenting biases in human judgements of other humans. One such bias is affinity bias, where recruiters prefer candidates with similar knowledge and skills as them (Bagues & Perez-Villadoniga, 2012). We show LM judges also systematically favor other models that make similar mistakes, generalizing previous results that showed LMs favor their own outputs (Liu et al., 2024; Panickssery et al., 2024). Overall, we believe AI evaluators should be accompanied with formal checks like consistency (Fluri et al., 2024).
@@
A second aspect of AI oversight is using another model’s supervision to train better models. This is similar to training on text generated by an LM (Chiang et al., 2023) with ongoing debates about its benefits (Kazdan et al., 2024), and an emerging paradigm of exploiting a gap in difficulty between solution generation and evaluation (Song et al., 2024). In this paper, we study the more established setup of training LMs on LM annotations, where Burns et al. (2024) demonstrated the phenomenon of weak to strong generalization, and it has been leveraged for other applications like image classification (Guo et al., 2024) and aligning models (Zhu et al., 2024). Prior work has attempted to understand weak to strong generalization, notably using “misfit error” (Charikar et al., 2024), which shows that the student’s disagreement with the weak supervisor after weak to strong training correlates with its accuracy gap from the weak supervisor. Instead, we show similarity between the weak supervisor and strong student can apriori predict gains from weak-to-strong training. The benefit of model diversity has previously been discussed in related settings like knowledge distillation for image classifiers (Roth et al., 2024) and training chess models that outperform the humans they are trained on (Zhang et al., 2024).

### v2

@@ Abstract
Model Similarity, AI Oversight, LLM as a Judge, Weak to Strong Generalization, Model Differences
@@ 1 Introduction
Machine Learning model capabilities have improved immensely over the last few years. Scaling up the amount of data used for training has played a crucial role in these improvements (Kaplan et al., 2020). Initially, most of the gains in Language Model (LM) capabilities came from scaling pretraining data (Llama Team, 2024a). Recently, there is increasing interest in post-training, either with human preferences (Ouyang et al., 2022), or task-specific expert annotations (Lightman et al., 2024). Collecting human preferences or annotations is slow and expensive. Therefore, with increasing model capabilities, an attractive alternative is to use LMs to annotate training data  (Gilardi et al., 2023) and score model outputs (Zheng et al., 2023), to boost both training (Stiennon et al., 2020) and evaluation (Li et al., 2024b). In this paper, we refer to both these techniques together as AI oversight [^1] The term is inspired by “scalable oversight” (Bowman et al., 2022), which studies human-in-the-loop mechanisms for AI Safety..
Can we rely on AI oversight going forward? This remains a topic of much debate. In this work, we study oversight from the perspective of model similarity. When assessing or teaching humans, it is well recognized that individuals have different strengths and weaknesses. Similarly, two models with 50% accuracy may misclassify completely different samples and thus be highly dissimilar (having different ‘strengths’). To measure model similarity, we build on error consistency (Geirhos et al., 2020), which measures overlap in the samples where two models err beyond what is expected by chance due to the two models’ accuracies. In Section 2, we extend the error consistency metric in two crucial ways – 1) by counting differences in predictions rather than correctness for each sample, and 2) incorporating output probabilities. In this way, our novel similarity metric, Chance Adjusted Probabilistic Alignment (CAPA), allows us to quantify functional similarity between models. We use this to analyze both evaluation and training using AI oversight as depicted in Figure 1:
@@ 2.2 Our Contribution
Figure 2: Metric comparison for independent models with uncorrelated predictions. In this simulation, for each model we select an independent random subset of samples as correct, with the first having a fixed $90\%$ accuracy, while for the second accuracy is varied from $50\%$ to $90\%$. CAPA correctly reports 0 similarity when models have uncorrelated errors.
[…]
@@ 3 Affinity Bias in LLM-as-a-Judge
Evaluating free-response model generations automatically at scale is tricky (Biderman et al., 2024). This has led to popular leaderboards like Arena-hard-auto (Li et al., 2024b), AlpacaEval 2.0 (Dubois et al., 2024), and AidanBench (McLaughlin et al., 2024) adopting LLM-as-a-Judge for scoring. Recent work cautions that language models are biased towards their own outputs (Liu et al., 2024; Panickssery et al., 2024), and these leaderboards assume that excluding the judge model from the rankings circumvents this problem. However, human interviewers have been shown to also be biased towards candidates with similar knowledge and traits, a phenomenon called affinity bias (Bagues & Perez-Villadoniga, 2012). We study whether LMs also exhibit a bias toward similar models. This would indicate that it is not sufficient to just use a held-out LM as the judge; one must account for the confounder of similarity.
@@ 3.1 Experimental Setup
Figure 3: Judgment Score Relation with Model Similarity on only across family pairs. Each line is a regression model fit between judgment and similarity scores. The circle shape indicates that only across-family judge-model pairs are plotted. We report for each fit the corresponding Pearson correlation values, $r$. We found significant positive correlation between judgment scores and CAPA across all judges, **∗ ∗ indicates $p<0.01$.
To study whether LM judges prefer more similar models, we evaluate a large set of judges and models on MMLU-Pro (Wang et al., 2024), a benchmark for hard problem solving questions across 14 domains. We filter 8,707 questions that can also be answered in a free-text response style, without options, following Myrzakhan et al. (2024). Each question is posed to every evaluated model as an MCQ and as an open-style question. The per-sample results of the former are used to compute the similarities of judge-model pairs, whereas responses to the latter are given to an LLM-as-a-judge. The judge has to make a binary decision on whether a free-text response is correct or wrong. This is done based on its own internal knowledge without access to a ground-truth solution. We call the average of binary judgments across the benchmark the model’s Judgment Score for a given judge. Using a parallel MCQ evaluation with ground-truth answers allows us to compare the judgment scores with verifiable accuracy measurements (details and comparisons with alternatives are in Appendix B.3), consistent with prior scalable oversight research (Bowman et al., 2022). We compute pairwise similarity with CAPA, $\kappa_{p}$, across 9 judge and 39 model pairs. For a complete overview of models investigated, question filtering, inference setup, and prompts used for judges, see Appendix B.
@@ 3.2 Results & Discussion
Q1: Do LM Judges favor more similar models? As a motivating example, Qwen2.5-72B-Instruct as a judge scores Qwen2.5-7B-Instruct as being 71% correct, while the more capable (41% vs 51% MCQ accuracy) Llama-3.1-70B-Instruct is deemed less accurate at 67%. In Figure 3 we show that model favoritism extends beyond self- or family- preference to all models that are functionally similar. We find a significant ($p<0.01$) positive correlation (average Pearson r=$0.84$) between LLM-as-a-judge scores and model similarity ($\kappa_{p}$) for all judges. To ensure this trend is not only an artifact caused by aggregating across all MMLU-Pro categories, per-category results for the same experiment are shown in the Appendix B.5. These exhibit the same affinity bias observed here for all different topics.
Q2: Is this merely due to better accuracy? Note that while $\kappa_{p}$ adjusts for inflation in agreement of highly accurate models, we do expect models with lower accuracy to be less similar with highly capable judge models, and thus have lower $\kappa_{p}$. To control for the accuracy of the evaluated model we perform multiple regression and partial correlation analysis (Table 2). The multiple regression analysis shows that both, accuracy and similarity, have a significant positive effect on the judge score. The coefficient of accuracy increases for more capable judge models, consistent with prior work showing improved alignment with human judges (Thakur et al., 2024). We find that especially for small models ($<$32B) the effect of similarity is greater. The partial correlation results control for accuracy and confirm that there is still a significant effect of similarity on judgment scores even for the best judge models. Altogether, the statistical analysis confirms that judgment scores are confounded by affinity bias. For extended results also including model size as a confounder please see Appendix B.1.
Q3: Does this transfer to LLM-as-a-judge for binary preference in free-form generation? In most cases, LLM-as-a-judge is not used to check the technical correctness of responses, but rather to match human preference by selecting the better one from a pair of responses. To extend our findings to the additional domain of chat responses, we provide an additional experiment on the AlpacaEval benchmark in Appendix B.4 (Dubois et al., 2023). We show that the Elo obtained by using the same judges as before to obtain binary preferences still strongly correlates with model similarity on MMLU-Pro. This indicates the existence of an affinity bias in settings other than QA.
@@ 4.1 Experimental Setup
Following Burns et al. (2024), we define “weak” and “strong” based on model size. We study 4 weak models in the $1-3$B parameter range, and 3 strong models in the $7-9$B parameter range, for a total of 12 model pairs, and $15$ of the binary classification NLP tasks studied in Burns et al. (2024), specified in Table 9. The smaller models do have lower accuracy (before finetuning) than the bigger ones across tasks, as shown in Appendix Figure 17, justifying calling them “weak” and “strong”. We measure similarity between the weak supervisor and base student model on the validation set. We then perform weak-to-strong training on the student model, using the confidence-weighted finetuning objective proposed in Burns et al. (2024). We investigate if similarity is an apriori predictor of performance gained on the test set. The full setup is consistent with the open-weight model reproduction by (Scherlis et al., 2024), and is described in Appendix C.1.
@@ 4.2 Results & Discussion
Q1: Does Complementary Knowledge Influence Performance Gain? Figure 4 shows that for all model combinations, similarity between the weak supervisor and initial strong student inversely correlates with the improvement obtained from weak-to-strong training ($r=-0.85)$. Even after using partial correlation analysis to control for the accuracy gap between the weak supervisor and strong student, similarity is inversely correlated with weak-to-strong gain ($r=-0.35,p<0.01$). Thus, tasks where the supervisor and student make less correlated errors tend to yield greater improvements. This contributes towards understanding why gains from weak to strong training vary across tasks, an open question posed by Burns et al. (2024). The observation is not specific to our similarity metric, Appendix Figure 16 shows the same trend holds with alternate metrics, albeit with lesser variance explained.
Q2. Does Complementary Knowledge Add Beyond Elicitation? The original explanation for performance gains from weak-to-strong generalization is that the weak supervisor “elicits” the latent knowledge in the superior representations of the stronger student (Burns et al., 2024). To investigate whether complementary knowledge adds to this explanation or is subsumed within it, we first obtain the strong model with “upper-bound” elicitation, by finetuning it on ground-truth annotations. We refer to this as the strong elicited model. We can then separate the test data into four parts based on whether the strong elicited and weak supervisor model were correct or wrong, measuring average accuracy of the weak-to-strong model on each part to disentangle gains from different factors. The experiment setup is discussed further in Appendix C.2.
@@
Q3. Where can weak-to-strong training improve? The strong elicited model is considered to represent upper-bound performance. However, recent work has shown that the performance of the strong elicited model can be surpassed (Shi et al., 2025), indicating it is not really an “upper-bound”. Indeed, the strong elicited model’s performance does not account for potential gains from leveraging the complementary knowledge of the weak supervisor. We compute a new ceiling, taking the union of correct predictions between the weak supervisor and strong elicited model, which is significantly higher as shown in Table 3. Interestingly, on the training set, the weak-to-strong trained model shows similar accuracy on the top-left and bottom-right quadrants as shown in Figure 15. Yet, when generalizing to unseen samples, it falls back more often to its initial priors. We hope this analysis guides future work on improving weak-to-strong training methodology, by highlighting leveraging complementary knowledge as a concrete avenue for improvement.
Figure 6: Average Similarity ($\kappa_{p}$) vs Model Capability. We split 130 LMs into 5 buckets based on their accuracy percentile. For each LM we compute its mean similarity within the bucket (across models from different developers), and plot it against model accuracy. The size of the scatter points indicates model size. As $\kappa_{p}$ measures overlap in mistakes, the positive correlation indicates LM mistakes are getting more correlated with increasing capabilities.
@@ 5.2 Results & Discussion
Q2. Why are model errors becoming more correlated? This is an interesting research direction in itself. We perform a preliminary analysis in Appendix D.2, summarizing key conclusions here. First, we observe only a slight increase in similarity for harder questions in our datasets, indicating difficulty is not a significant confounder for this trend. We find this trend is stronger in instruction-tuned models, and using alternative architectures like Mamba (Gu & Dao, 2024) may not be enough to increase diversity.
@@ 6 Related Work
There is increasing interest in finding differences between models for applications like visual tools for comparative analytics (Strobelt et al., 2021; Kahng et al., 2024), efficient human evaluation (Boubdir et al., 2023), comparing learning algorithms (Shah et al., 2023), identifying side-effects of API updates (Eyuboglu et al., 2024) or quantization (Dutta et al., 2024). Prior work has also looked at qualitatively describing differences between data distributions (Zhong et al., 2022, 2023; Dunlap et al., 2024, 2025). Our work proposes metrics to quantify LM differences (or similarity). Huh et al. (2024) used representation similarity metrics (Kornblith et al., 2019; Bansal et al., 2021) to show convergence in visual representations and their alignment with language representations. In contrast, we show model mistakes are becoming more correlated as capabilities improve, using sample level evaluations (Burnell et al., 2023) such as those available on OpenLLMLeaderboard (Myrzakhan et al., 2024) and HELM (Bommasani et al., 2023)Geirhos et al. (2020) proposed measuring “error consistency” between image classifiers and humans, with Geirhos et al. (2021) showing an early trend of data-rich models making more similar mistakes to humans. We enrich this metric, distinguishing between different mistakes and incorporating probabilistic information. Bommasani et al. (2023) used a similar metric to demonstrate homogenization in outcomes when using LMs for decision making. Our work shows that such risks from algorithmic monoculture (Kleinberg & Raghavan, 2021) are increasingly relevant as LM capabilities improve.
@@
Our results on AI judges fall in a broader line highlighting their pitfalls (Zheng et al., 2025). These include biases such as favoring verbose texts or options at certain positions (Koo et al., 2024; Ye et al., 2025). Interestingly, these biases are also sometimes found in human annotators (Chen et al., 2024). In fact, there is rich literature documenting biases in human judgements of other humans. One such bias is affinity bias, where recruiters prefer candidates with similar knowledge and skills as them (Bagues & Perez-Villadoniga, 2012). We show LM judges also systematically favor other models that make similar mistakes, generalizing previous results that showed LMs favor their own outputs (Liu et al., 2024; Panickssery et al., 2024). Overall, we believe AI evaluators should be accompanied with formal checks like consistency (Fluri et al., 2024).
@@
A second aspect of AI oversight is using another model’s supervision to train better models. This is similar to training on text generated by an LM (Chiang et al., 2023) with ongoing debates about its benefits (Kazdan et al., 2025), and an emerging paradigm of exploiting a gap in difficulty between solution generation and evaluation (Song et al., 2025). In this paper, we study the more established setup of training LMs on LM annotations, where Burns et al. (2024) demonstrated the phenomenon of weak to strong generalization, and it has been leveraged for other applications like image classification (Guo et al., 2024) and aligning models (Zhu et al., 2025). Prior work has attempted to understand weak to strong generalization, notably using “misfit error” (Charikar et al., 2024), which shows that the student’s disagreement with the weak supervisor after weak to strong training correlates with its accuracy gap from the weak supervisor. Instead, we show similarity between the weak supervisor and strong student can apriori predict gains from weak-to-strong training. The benefit of model diversity has previously been discussed in related settings like knowledge distillation for image classifiers (Roth et al., 2024) and training chess models that outperform the humans they are trained on (Zhang et al., 2024).